# with Language Model Diagnostic Assistance

### v1

@@ 4 ED-Copilot for Diagnostic Assistance
With MIMIC-ED-Assist, we develop ED-Copilot, a method to provide (time)-cost-effective diagnostic assistance for ordering informative laboratory groups to flag patients at high risk for critical outcome and lengthened ED stay. We first provide a high-level description of ED-Copilot, followed by how it is trained. Training ED-Copilot involves a supervised fine-tuning stage for adapting the pre-trained language model (PLM) (e.g. BioGPT) backbone to our task, followed by an RL stage that selects laboratory groups to reduce time-cost. At the end of this section, we discuss how ED-Copilot conducts inference.
Figure 1: Overview of ED-Copilot training flow on one ED visit.
@@ 4.1 Problem Formulation
Task. The task for ED-Copilot is to assist doctors by recommending the next laboratory group to flag patients with potential critical outcome or lengthened ED stay, based on patient triage and laboratory information.
Let a patient have one triage group $x_{0}$ and $n$ laboratory groups $[x_{i}]_{i=1}^{n}$with their respective associated results $[r_{i}]_{i=0}^{n}$ in the order of $[x_{0},r_{0},x_{1},r_{1},x_{2},r_{2},....,x_{n},r_{n}]$, along with a diagnostic outcome $y$. ED-Copilot first linearizes this patient’s triage and laboratory into a text sequence, then passes it into a language model backbone to suggest the next lab group or predict the outcome.
@@
Language Model Backbone. ED-Copilot uses a PLM $G_{\theta}$ to take the input sequence, append a [EOS] token after each group, and uses classification head upon the hidden representations of $G_{\theta}$ to select next laboratory group or predict the outcome:
@@
Here $G_{\theta}$ generates hidden representations for each token, on which we use a multi-layer perceptron (MLP) $p_{\phi}(x_{i}|\textbf{{h}}_{<i})$ to predict the next laboratory group from the laboratory label space and another MLP $p_{\psi}(y|\textbf{{h}}_{\leq n})$ to predict the diagnostic outcome $y$, conditioned on the hidden representation h of the previous [EOS] position. Note that we only predict laboratory groups, not their associated laboratory tests’ results, which should be filled by running the actual lab testing.
@@ 4.2 Supervised Fine-tuning
The first stage of training uses standard language model fine-tuning techniques, where we train the PLM to suggest the next laboratory group to order and eventually to predict target outcomes. To predict the next laboratory group, we use a standard auto-regressive loss function:
@@
For outcome prediction, we use the following loss
@@
Then, we fine-tune the PLM by:
$\min_{\theta,\psi,\phi}(\mathcal{L}_{\text{lab}}+\mathcal{L}_{y}),$ (6)
@@ 4.3 Reinforcement Learning
Choosing informative laboratory groups is a key factor in reducing laboratory testing, and hence improving wait-time, and reducing ED LOS. In this section, we employ RL to introduce the notion of time-cost effectiveness to the supervised fine-tuned PLM, to select laboratory groups that maximize F1-score while minimizing time-cost.
@@
Action Space. The action space $\mathcal{A}=\{x_{1},x_{2},....,x_{K}\}\cup\{y^{+},y^{-}\}$ include $K$ number of groups, with action $a\in\{x_{1},x_{2},....,x_{K}\}$ corresponding to ordering a laboratory test group, with a time-cost $c(a)$. Action $a\in\{y^{+},y^{-}\}$ means predicting the outcome and terminating the process.
Policy. The policy $\pi_{\eta}:\mathcal{S}\rightarrow\mathcal{A}$ maps from states to actions, and is parameterized by weights $\eta$. Specifically, $\pi_{\eta}(a|\textbf{{h}}_{<i})$ outputs probability of an action $a$ using the hidden representation $\textbf{{h}}_{<i}$ corresponding to [EOS] token for state $s_{i}$.
RL Training. We train the policy $\pi_{\eta}$ to follow two objectives: maximize F1-score and minimize time-cost. We measure the time cost of policy $\pi_{\eta}$, as follows:
@@
where $\mathbb{E}_{\pi_{\eta}}$ is the expectation under policy $\pi_{\eta}$. Using this definition of cost, let $\pi^{*}_{\eta}(\alpha,\beta)$ represent the policy that maximizes F1-score while minimizing time-cost for hyper-parameters $\alpha,\beta$ to be defined. Then, $\pi_{\alpha,\beta}^{*}$ is equivalent to the solution of the following program:
$\pi^{*}_{\eta}(\alpha,\beta)=\text{argmax}_{\pi_{\eta}}\{\text{TN}(\pi_{\eta})+\alpha\text{TP}(\pi_{\eta})+\beta\text{Cost}(\pi_{\eta})\}.$ (8)
TN$(\pi_{\eta})$ and TP$(\pi_{\eta})$ are the numbers of true negatives and true positives, under policy $\pi_{\eta}$. Hyper-parameters $\alpha,\beta$ control the trade-off between F1-score and time-cost. We train the MLP via proximal policy optimization (Schulman et al., 2017).
Measuring time-cost. We measure time-cost via the total time taken to run all laboratory groups. In the ED, laboratory groups are often ordered both in parallel, and sequentially since the decision on new tests to order depends on previous tests. Additionally, the number of tests an ED clinician orders also depends on other factors outside of the patient’s health record, such as insurance policy. As the first step towards reducing ED LOS with AI, this work uses the sum of test time as the optimization target, which is an approximation of the influence of test costs on total ED LOS. Better ways to model ED LOS are future research for both the AI and healthcare communities.
Table 2: Overall performances of ED-Copilot and baselines in predicting Critical Outcome (transfer to ICU or mortality) and Lengthened ED Stay (spending more than 24 hours in emergency room), and the Average Time Cost (in minutes) to make a prediction. Sensitivity and specificity are the true positive and true negative rates.
@@
Model F1 AUC Sensitivity Specificity Avg. Time Cost F1 AUC Sensitivity Specificity Avg. Time Cost
Random Forest 0.377 0.807 0.754 0.748 265 Minutes 0.206 0.698 0.693 0.616 265 Minutes
XGBoost 0.379 0.807 0.731 0.744 265 Minutes 0.212 0.679 0.619 0.661 265 Minutes
LightGBM 0.394 0.813 0.725 0.769 265 Minutes 0.217 0.705 0.706 0.605 265 Minutes
3-layer DNN 0.339 0.743 0.676 0.683 265 Minutes 0.194 0.637 0.649 0.593 265 Minutes
SM-DDPO 0.353 0.780 0.685 0.763 182 Minutes 0.183 0.619 0.472 0.739 177 Minutes
ED-Copilot 0.413 0.820 0.750 0.779 125 Minutes 0.252 0.707 0.725 0.606 154 Minutes
@@ 4.4 Inference
During inference, ED-Copilot assists ED clinicians to optimize their workflow by suggesting the next most informative laboratory group to order, and also by flagging patients at high risk of critical outcome and lengthened ED stay. Specifically, given triage information and previous test results, ED-Copilot recommends additional tests. The results for these additional tests can then be fed back to ED-Copilot to suggest additional laboratory groups or to flag one of two possible outcomes. Further, ED-Copilot can be used in more flexible ways depending on a clinician’s needs. For example, multiple tests can be ordered using multiple suggestions from ED-Copilot.
@@
Figure 2: Prediction accuracy and average number of laboratory groups of ED-Copilot with different maximum allowed times to perform lab tests. Each point reflects ED-Copilot’s F1/AUC (y-axes) at different time thresholds.
@@ 5 Experimental Set-up
Dataset Split. We randomly split the dataset using $80\%$ for training, $10\%$ for validation, and $10\%$ for testing, while ensuring each split has the same class distribution. The validation set is used to tune hyper-parameters for all methods. During inference, the initial state of that patient is set to be demographic and triage information.
Evaluation Metrics. We use four evaluation metrics: F1-score, area under the receiver operating characteristic (AUC), sensitivity (true positive rate), and specificity (true negative rate), all standard in healthcare tasks prediction tasks (Harutyunyan et al., 2019; Xie et al., 2022).
Baselines. For prediction tasks only, ED-Copilot is compared to three tree-based methods: random forests (Breiman, 2001), XGBoost (Chen & Guestrin, 2016) and LightGBM (Ke et al., 2017), all of which are known to achieve strong performance on tabular data. Additionally, we compare ED-Copilot to a $3$-layer dense deep neural network (DNN).
Note that since each patient is not observed under all lab groups, there are many missing values in our dataset. For these baselines, we consider a number of imputation methods: (1) mean imputation, (2) median imputation, and (3) zero-imputation (i.e., replacing missing values by 00), (4) using a dummy indicator to encode missing values. Results with the best imputation method is presented in the main manuscript, and the rest are in Appendix C.
We also compare ED-Copilot to another cost-effective baseline, SM-DDPO (Yu et al., 2023), which selects laboratory groups to minimize time cost while maximizing F1-score. SM-DDPO is a non-personalized method, and selects the same laboratory groups for all patients. All hyper-parameters are tuned on the validation set to maximize F1-score.
Time Cost. Each group of lab tests is assigned a time cost by observing time-stamps in the MIMIC-IV database. For both ED-Copilot, and all baselines, we measure the amount of time taken by averaging the time required for an algorithm to make a prediction across all patients. This reflects the impact of laboratory tests on ED LOS.
Implementation Details. BioGPT (Luo et al., 2022) is used as our backbone language model, which is a generative pre-trained transformer (347 million parameters) for biomedical text generation and mining. In the RL phase of training ED-Copilot, for a given patient, our action space is restricted to laboratory groups that the patient actually performed. A discussion when selecting unobserved laboratory groups can be found in Section 6.4.
@@
F1 AUC Sensitivity Specificity Avg. Time Cost F1 AUC Sensitivity Specificity Avg. Time Cost
ED-Copilot (345M) 0.413 0.820 0.750 0.779 125 Minutes 0.252 0.707 0.725 0.606 154 Minutes
Linearization Raw Lab Test Name 0.397 0.777 0.768 0.677 134 Minutes 0.241 0.695 0.611 0.701 144 Minutes
@@
w/o. Chemical 0.420 0.827 0.788 0.746 0.234 0.702 0.656 0.606
Backbone Pythia (70M) 0.290 0.698 0.574 0.702 166 Minutes 0.178 0.596 0.555 0.619 126 Minutes
GPT2-Medium (345M) 0.358 0.757 0.621 0.747 133 Minutes 0.166 0.539 0.498 0.584 96 Minutes
@@ 6 Evaluation Results
6.1 Prediction Accuracy and Time Cost
Overall Performance. The prediction performance and time costs of all methods are presented in Table 2. For critical outcome, ED-Copilot outperforms the next best baseline by 1.9%, 0.7%, 2.5%, and 1% for F1-score, AUC, sensitivity, and specificity, respectively. For lengthened ED stay, ED-Copilot outperforms the next best model by 3.5%, 0.2%, 1.9%, and 0.1% for the four metrics. ED-Copilot halves average time cost from roughly $4$ hours to $2$ hours.
In comparison to the other cost-effective baseline, SM-DDPO, ED-Copilot also has significantly better accuracy and lower time cost. Both methods utilize the same RL algorithm to minimize time-cost, but ED-Copilot benefits from the strong capability of PLM backbone (Sec 6.2) and the personalized assistance it enabled (Sec 6.3).
Prediction Accuracy at Different Time Costs. In Figure 2, we investigate how ED-Copilot’s prediction performance changes as we vary the maximum time allowed to perform laboratory tests. The prediction performance increases with the maximum allowed time and number of laboratory groups selected. ED-Copilot requires relatively few tests to approximate peak performance for lengthened ED stay, while prediction accuracy for critical outcome steadily increases with more laboratory tests allowed. These curves illustrate different trade-offs of diagnostic assistance for different outcomes and help ED clinicians better leverage when allocating resources.

### v2

@@ 4 ED-Copilot for Diagnostic Assistance
We use MIMIC-ED-Assist to develop ED-Copilot, which offers cost-effective diagnostic assistance by minimizing the number of laboratory groups required to identify high-risk patients. This section is organized as follows. First, we provide a high-level overview of ED-Copilot, see Figure 1 for a visualization. Then, we detail its training process which consists of two stages: supervised fine-tuning to adapt the pre-trained language model (PLM) (e.g., BioGPT) to our prediction task, followed by RL to select laboratory groups that reduce time-costs. Finally, we discuss how ED-Copilot conducts inference.
Figure 1: Overview of ED-Copilot procedure on one ED visit.
@@ 4.1 Problem Formulation
Let a patient has triage features denoted by $x_{0}$ and $n$ laboratory groups $[x_{i}]_{i=1}^{n}$with their associated results $[r_{i}]_{i=0}^{n}$ in order $[x_{0},r_{0},x_{1},r_{1},x_{2},r_{2},....,x_{n},r_{n}]$. Additionally, let $y$ denotes our prediction targets defined in Section 3. ED-Copilot first linearizes this patient’s triage and laboratory results into a text sequence, which is then inputted into the PLM backbone to suggest the next group of tests or predict the outcome. We describe these steps in detail as follows.
@@
Language Model Backbone. We apply PLM $G_{\theta}$ to the text sequence constructed above to obtain hidden representations H for each [EOS] token. We then use a MLP $p_{\phi}(x_{i}|\textbf{{h}}_{<i})$ to predict the next laboratory group using hidden representations $\textbf{{h}}_{<i}$ for tokens [EOS]<i. Lastly, we predict outcomes $y$ using $p_{\psi}(y|\textit{{h}}_{\leq n})$ where $\textit{{h}}_{n}$ is the hidden state for token [EOS]n. This process is reflected below.
@@
Note that we only predict laboratory groups, not their associated laboratory tests’ results, which should be determined by conducting the actual laboratory tests. Next, we describe our training procedure.
@@ 4.2 Supervised Fine-tuning
In this section, we perform supervised fine-tuning of our PLM to suggest the next laboratory group and predict outcomes. To predict the next laboratory group, we use a standard auto-regressive loss function,
@@
For outcome prediction, we use the following loss,
@@
To fine-tune, we minimize the loss,
$\theta^{*},\psi^{*},\phi^{*}=\min_{\theta,\psi,\phi}(\mathcal{L}_{\text{lab}}+\mathcal{L}_{y}),$ (6)
@@ 4.3 Reinforcement Learning
Choosing informative laboratory groups is a key factor in reducing laboratory testing, and thereby reducing ED LOS. In this section, we employ RL to introduce the notion of time-cost effectiveness to the fine-tuned PLM to select laboratory groups that maximize predictive accuracy while minimizing time-cost.
@@
Action Space. Denote the action space as $\mathcal{A}=\{x_{1},x_{2},....,x_{K}\}\cup\{y^{+},y^{-}\}$. Action $a\in\{x_{1},x_{2},....,x_{K}\}$ corresponding to ordering a group of laboratory tests with associated time-cost $c(a)$. Action $a\in\{y^{+},y^{-}\}$ refers to predicting an outcome and terminating the MDP.
Policy. The policy $\pi_{\eta}:\mathcal{S}\rightarrow\mathcal{A}$ maps from states to actions, and is parameterized by $\eta$. Specifically, $\pi_{\eta}(a|\textbf{{h}}_{<i})$ outputs probability of an action $a$ using the hidden representation $\textbf{{h}}_{<i}$ corresponding to [EOS] token for state $s_{i}$.
RL Training. We train the policy $\pi_{\eta}$ to follow two objectives: maximize F1-score and minimize time-cost. We measure the time-cost of policy $\pi_{\eta}$, as follows:
@@
where $\mathbb{E}_{\pi_{\eta}}$ is the expectation under policy $\pi_{\eta}$. Using (7), let $\pi^{*}_{\eta}(\alpha,\beta)$ represent the policy that maximizes F1-score while minimizing time-cost for hyper-parameters $\alpha,\beta$ to be defined. Then, $\pi_{\alpha,\beta}^{*}$ is equivalent to the solution of the following program:
$\begin{split}\pi^{*}_{\eta}(\alpha,\beta)=\text{argmax}_{\pi_{\eta}}\{&\text{TN}(\pi_{\eta})+\alpha\text{TP}(\pi_{\eta})\\ &+\beta\text{Cost}(\pi_{\eta})\}.\end{split}$ (8)
TN$(\pi_{\eta})$ and TP$(\pi_{\eta})$ are the numbers of true negatives and true positives under policy $\pi_{\eta}$. Hyper-parameters $\alpha,\beta$ control the trade-off between F1-score and time-cost. We train the MLP via proximal policy optimization (Schulman et al., 2017).
Measuring Time-cost. We measure time-cost via the total time taken to run all laboratory groups. In the ED, laboratory groups are often ordered both in parallel, and sequentially since the decision on new tests to order depends on previous tests. Additionally, the number of tests an ED clinician orders depends on other factors outside of the patient’s health record, such as insurance policy. As the first step towards reducing ED LOS with AI, we use the sum of laboratory group time-costs as the total cost, which serves as an approximation to the effect of testing on ED LOS. Better ways to model ED LOS are future research for both the AI and healthcare communities.
Table 2: Overall performance of ED-Copilot and baselines in predicting Critical Outcome (transfer to ICU or mortality) and Lengthened ED Stay (exceeding 24 hours in the ED), as well as the Average Time-cost (in minutes) to perform laboratory tests. Sensitivity and specificity are true positive and true negative rates. We report results averaged over three random seeds alongside standard deviations.
@@
Model F1 AUC Sensitivity Specificity Avg. Time-cost F1 AUC Sensitivity Specificity Avg. Time-cost
Random Forest 0.377 (0.015) 0.807 (0.011) 0.754 (0.012) 0.748 (0.005) 265 Min 0.206 (0.014) 0.698 (0.011) 0.693 (0.016) 0.616 (0.024) 265 Min
XGBoost 0.379 (0.019) 0.807 (0.009) 0.731 (0.017) 0.744 (0.006) 265 Min 0.212 (0.010) 0.679 (0.007) 0.619 (0.020) 0.661 (0.020) 265 Min
LightGBM 0.394 (0.016) 0.813 (0.008) 0.725 (0.012) 0.769 (0.004) 265 Min 0.217 (0.015) 0.705 (0.011) 0.706 (0.017) 0.605 (0.014) 265 Min
3-layer DNN 0.339 (0.032) 0.743 (0.021) 0.676 (0.024) 0.683 (0.011) 265 Min 0.194 (0.031) 0.637 (0.013) 0.649 (0.015) 0.593 (0.014) 265 Min
SM-DDPO 0.353 (0.031) 0.780 (0.020) 0.685 (0.023) 0.763 (0.022) 182 (32) Min 0.183 (0.028) 0.619 (0.012) 0.472 (0.012) 0.739 (0.011) 177 (60) Min
ED-Copilot 0.413 (0.028) 0.820 (0.021) 0.750 (0.018) 0.779 (0.011) 125 (21) Min 0.232 (0.023) 0.707 (0.015) 0.725 (0.018) 0.606 (0.015) 154 (33) Min
@@ 4.4 Inference
During inference, ED-Copilot assists clinicians to optimize their workflow by suggesting the next most informative laboratory group, and also by flagging patients at high risks of critical outcome and lengthened ED stay. Specifically, given triage information and previous test results, ED-Copilot recommends additional tests. The results of these tests are fed back to ED-Copilot to suggest additional laboratory groups or to flag one of two possible outcomes. Further, ED-Copilot can be used in more others ways depending on clinical needs. For example, multiple tests can be ordered using multiple suggestions from ED-Copilot.
@@
Figure 2: Prediction accuracy and average number of laboratory groups of ED-Copilot with different maximum allowed time to perform laboratory tests. Each point reflects ED-Copilot’s F1/AUC (y-axes) at different time upper-bounds.
@@ 5 Experimental Set-up
Dataset Split. We randomly split the dataset using $80\%$ for training, $10\%$ for validation, and $10\%$ for testing, while ensuring each split has the same class distribution. The validation set is used to tune hyper-parameters. During inference, the initial state of a patient is set to their triage information.
Evaluation Metrics. We use four evaluation metrics: F1-score, area under the receiver operating characteristic (AUC), sensitivity (true positive rate), and specificity (true negative rate) which are all standard in healthcare tasks prediction tasks (Harutyunyan et al., 2019; Xie et al., 2022).
Baselines. For prediction tasks only, ED-Copilot is compared to three tree-based methods: random forests (Breiman, 2001), XGBoost (Chen & Guestrin, 2016) and LightGBM (Ke et al., 2017), all of which are known to achieve strong performance on tabular data. We also compare ED-Copilot to a $3$-layer dense deep neural network (DNN).
Since each patient is not observed under all laboratory groups, there are many missing values in MIMIC-ED-Assist. Since baselines cannot naturally handle missing values, we consider a variety of imputation methods: (1) mean imputation, (2) median imputation, and (3) zero-imputation (i.e., replacing missing values by 00), (4) using a dummy indicator to encode missing values. Results for the best imputation method are presented in the main manuscript. The rest are in Appendix C.
We compare ED-Copilot to another cost-effective baseline, SM-DDPO (Yu et al., 2023), which selects laboratory groups to minimize time-cost while maximizing F1-score. SM-DDPO is a non-personalized method, i.e., it selects the same laboratory groups for all patients.
Time-cost. As discussed, each group of tests is assigned a time-cost by observing time-stamps in the MIMIC-IV database. For all methods, we measure the amount of time taken by averaging the time required for an algorithm to make a prediction across patients.
Implementation Details. BioGPT (Luo et al., 2022) is used as our backbone language model, which is a generative pre-trained transformer (347 million parameters) for biomedical text generation and representation. In the RL phase of training ED-Copilot, for a given patient, our action space is restricted to laboratory groups that the patient actually performed. Experiments when ED-Copilot is allowed to select unobserved laboratory groups can be found in Section 6.5.
@@
F1 AUC Sensitivity Specificity Avg. Time-cost F1 AUC Sensitivity Specificity Avg. Time-cost
ED-Copilot (345M) 0.413 0.820 0.750 0.779 125 Min 0.252 0.707 0.725 0.606 154 Min
Linearization Raw Lab Test Name 0.397 0.777 0.768 0.677 134 Min 0.241 0.695 0.611 0.701 144 Min
@@
w/o. CHEM 0.420 0.827 0.788 0.746 0.234 0.702 0.656 0.606
Backbone BioGPT (345M) w/o. RL 0.381 0.810 0.725 0.765 265 Min 0.236 0.718 0.710 0.620 265 Min
Llama (7B LORA) w/o. RL 0.397 0.798 0.692 0.767 265 Min 0.232 0.701 0.705 0.610 265 Min
[…]
GPT2-Medium (345M) w. RL 0.358 0.757 0.621 0.747 133 Min 0.166 0.539 0.498 0.584 96 Min
@@ 6 Evaluation Results
In this section, we present our experimental results. Section 6.1 compares the predictive and time-cost performance of ED-Copilot with baselines. Section 6.2 describes our ablation studies. Sections 6.3 and 6.4 establish ED-Copilot’s ability to personalize recommendations based on patient severity and achieve consistent performance across subgroups such as age and sex. Lastly, in Section 6.5, we investigate ED-Copilot’s performance when it is not restricted to select administered groups.
6.1 Prediction Accuracy and Time-cost
Overall Performance. The prediction performance and time-costs of all methods are presented in Table 2. For critical outcome, ED-Copilot outperforms the next best baseline by 1.9%, 0.7%, 2.5%, and 1% for F1-score, AUC, sensitivity, and specificity, respectively. For lengthened ED stay, ED-Copilot outperforms the next best model by 3.5%, 0.2%, 1.9%, and 0.1% for the four metrics. ED-Copilot halves average time-costs from roughly $4$ hours to $2$ hours.
In comparison to the other cost-effective baseline, SM-DDPO, ED-Copilot also has significantly better accuracy and lower time-costs. Both methods utilize the same RL algorithm to minimize time-cost, but ED-Copilot benefits from the strong capability of PLM backbone (Sec 6.2) and the personalized diagnostic assistance it enabled (Sec 6.3).
Prediction Accuracy at Different Time-costs. In Figure 2, we investigate how ED-Copilot’s prediction performance changes as we vary the maximum time allowed to perform laboratory tests. Performance increases with maximum allowed time and number of laboratory groups selected. ED-Copilot requires relatively few tests to achieve peak performance for lengthened ED stay, while prediction accuracy for critical outcome steadily increases as more tests are allowed. These curves illustrate different requirements for different outcomes and shed light as to how clinicians should allocate resources across tasks.
6.2 Ablation Studies
Linearization Technique. Following Hegselmann et al. (2023) and Manikandan et al. (2023), we replace the true laboratory test name with a standard feature name while keeping the result for that laboratory group unchanged. That is, we linearize laboratory results as follows ( feature1 : value1 |||| feature2 : value2 … ). Table 3 shows using standard feature names leads to better F1 and AUC score as compared to using the raw laboratory test names. One possible reason is that using raw laboratory test names biases ED-Copilot to pick names over informativeness for prediction.
Feature Importance. We examine the change in ED-Copilot performance when training with and without a given feature set from top three popular feature sets: triage information, complete blood count (CBC), and chemical group (CHEM). Table 3 shows removing triage features leads to a significant drop in performance. Our ED collaborators confirms this is medically plausible since initial laboratory tests are based on triage information. Our results also show all laboratory groups are not equally important, for example, removing CBC causes a larger drop in performance than removing the chemical laboratory group.